OpenStudy (anonymous):

A sailor pulls a boat horizontally alongside a dock for 10.0 m with an angled rope. He does 766 J of work and has a force of 100.0 N. What was the angle of the rope? (Points : 3) 30° 40° 50° 60°

jimthompson5910 (jim_thompson5910):

\[\Large w = f*d*cos(\theta)\] \[\Large 766 = 100*10*cos(\theta)\] solve for theta

OpenStudy (anonymous):

i got .17 is that right? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

ops sorry wrong thing i got 60 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

try solving 766 = 100*10*x for x

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

I got 1.3 but i think i did it wrong

jimthompson5910 (jim_thompson5910):

yeah you did

OpenStudy (anonymous):

can you explain

jimthompson5910 (jim_thompson5910):

does this help 766 = 100*10*x 766 = 1000x ?

OpenStudy (anonymous):

then dont you divide by 766?

jimthompson5910 (jim_thompson5910):

you divide both sides by 1000 to move the 1000 over

OpenStudy (anonymous):

then you get .766

jimthompson5910 (jim_thompson5910):

cos(theta) = 0.766 theta = ???

OpenStudy (anonymous):

Im not sure how to solve for that

jimthompson5910 (jim_thompson5910):

apply arccosine to both sides

OpenStudy (anonymous):

and i do that by adding x on both sides?

jimthompson5910 (jim_thompson5910):

what is the arccosine of 0.766?